# DeepTopoNet: A Framework for Subglacial Topography Estimation on the Greenland Ice Sheets

### v1

@@ Abstract.
Understanding Greenland’s subglacial topography is critical for projecting the future mass loss of the ice sheet and its contribution to global sea-level rise. However, the complex and sparse nature of observational data, particularly information about the bed topography under the ice sheet, significantly increases the uncertainty in model projections. Bed topography is traditionally measured by airborne ice-penetrating radar that measures the ice thickness directly underneath the aircraft, leaving data gap of tens of kilometers in between flight lines. This study introduces a deep learning framework, which we call as DeepTopoNet, that integrates radar-derived ice thickness observations and BedMachine Greenland data through a novel dynamic loss-balancing mechanism. Among all efforts to reconstruct bed topography, BedMachine has emerged as one of the most widely used datasets, combining mass conservation principles and ice thickness measurements to generate high-resolution bed elevation estimates. The proposed loss function adaptively adjusts the weighting between radar and BedMachine data, ensuring robustness in areas with limited radar coverage while leveraging the high spatial resolution of BedMachine predictions i.e. bed estimates. Our approach incorporates gradient-based and trend surface features to enhance model performance and utilizes a convolutional neural network (CNN) architecture (i.e., BedTopoCNN) designed for subgrid-scale predictions. By systematically testing on the Upernavik Isstrøm) region, the model achieves high accuracy (MAE: 12.49, RMSE: 19.38, and R2: 0.99), outperforming baseline methods in reconstructing subglacial terrain. This work demonstrates the potential of deep learning in bridging observational gaps, providing a scalable and efficient solution to inferring subglacial topography. The framework offers implications for ice sheet modeling, paving the way for improved predictions of ice sheet flow and their impact on sea level rise.
@@ 1. Introduction
The Greenland Ice Sheet is a major contributor to global sea-level rise, driven by dynamic processes such as ice flow and iceberg calving. Rising sea levels pose a critical threat to coastal communities worldwide, increasing the risk of flooding, shoreline erosion, and habitat loss (Team et al., 2023). Estimating these contributions requires a detailed understanding of subglacial topography, which influences ice flow and basal friction (Barnes et al., 2021; Åkesson et al., 2021). However, accurately inferring the bed topography remains a significant challenge due to limited observational coverage, especially in the ice sheet interior. Traditional numerical methods, such as mass conservation models and geostatistic techniques (Morlighem et al., 2011; Bamber et al., 2001, 2013), have been instrumental but often struggle with sparse data.
@@
To summarize, the contributions of this paper are threefold: (i) DeepTopoNet Framework: A novel deep learning framework integrating radar observations and BedMachine data, featuring a CNN architecture enhanced with gradient-based covariates and trend-surface features; (ii) Dynamic Loss-Balancing Mechanism: An adaptive mechanism ensuring robust performance in regions with sparse radar data by balancing contributions from multimodal data sources; and (iii) Comprehensive Evaluation: Extensive benchmarking across Greenland’s Upernavik and unseen regions, demonstrating superior performance over traditional interpolation and machine learning methods.
@@ 2. Related Work
Subglacial Bed Topography Prediction. Inferring the subglacial bed topography has been a critical focus in glaciology due to its implications for ice sheet mass balance and their contributions to global sea-level rise. Recent advancements have integrated machine learning, physics-based models, and hybrid approaches to tackle this complex problem. Cheng et al. (2024) (Cheng et al., 2024) applied Physics-Informed Neural Networks (PINNs) to model ice flow dynamics for Helheim Glacier in Greenland. The study demonstrated PINNs’ capability to infer basal friction coefficients and interpolate sparse ice thickness observations with high accuracy. The framework uniquely incorporates the conservation of momentum (i.e., the shallow-shelf/shelfy-stream approximation) (MacAyeal, 1989; Morland, 1987) as soft constraints in the loss function, improving predictions where data are sparse and enhancing the versatility of the model. Yi et al. (2023) (Yi et al., 2023) evaluated nine machine learning and statistical models, including XGBoost, Gaussian Process Regression, and hybrid models, to predict Greenland’s subglacial bed topography. The study highlighted the strong performance of XGBoost when paired with kriging-based interpolation, achieving the best RMSE and R2. Similarly, Leong and Horgan (2020) (Leong and Horgan, 2020) introduced DeepBedMap, a deep learning framework based on generative adversarial networks, to produce high-resolution bed elevation maps of Antarctica. Their approach combined multi-resolution spatial data, such as surface elevation and ice velocity, to improve the representation of subglacial terrain roughness. These studies underscore the importance of combining domain knowledge with advanced computational tools to address challenges in regions with sparse observations. This paper builds on these foundations, proposing a robust framework that integrates multimodal data, deep learning, and loss balancing to further enhance predictive performance.
@@ 4.1. Prior Physical Model (BedMachine)
BedMachine (Morlighem et al., 2017, 2020) is the most widely used subglacial topography map of Greenland. It employs a mass conservation approach in fast flowing regions, combining radar-derived ice thickness measurements, satellite velocity data, and surface mass balance reconstructions to produce a high-resolution bed topography map. The idea of mass conservation is to infer the bed elevation by ensuring conservation of mass: , where is the ice velocity, is surface elevation, is bed elevation, and is the apparent mass balance that combines observed thickness change and surface mass balance from a regional climate model. BedMachine ensures physically realistic reconstructions in regions of fast-flowing ice, but remains highly uncertain in slower moving regions where it is based on a simple interpolation scheme that connects existing radar data. The output is a seamless topographic map with horizontal resolution up to 150 m. In this study, BedMachine data serves as a reference for the target variable, subglacial bed topography (). The proposed DeepTopoNet framework goes beyond the interpolation based approach and the reliance on mass conservation.
@@ 5.1.1. Dataset
The dataset used in this study is obtained from the Upernavik Isstrøm region in the Central West of Greenland, encompassing four distinct sub-regions. Each sub-region covers a spatial extent of grid cells, representing an area with diverse ice dynamics and varying data density. The number of radar data points differs across the sub-regions, reflecting the uneven distribution of observations. The dataset includes several observed variables critical for subglacial topography inference:
@@ 5.3. Quantitative Comparison
Compared to traditional interpolation methods, such as IDW (Johnston et al., 2001) and RBF (Heryudono and Driscoll, 2010), DeepTopoNet delivers significantly improved performance. For instance, while IDW achieves moderate results in some sub-regions (e.g., SSIM of 0.87 in Sub-region II), its RMSE values remain substantially higher (e.g., 101.05 in Sub-region I). Similarly, RBF fails to capture finer spatial details, resulting in higher errors across all metrics. In contrast, DeepTopoNet achieves considerably lower RMSE and higher SSIM, demonstrating its ability to model spatial variability more effectively than interpolation methods.
Against pixel-based machine learning methods, including RF (Sekulić et al., 2020), MLP-PE (Li et al., 2021) and VAE+XGBoost (Yi et al., 2023), DeepTopoNet exhibits a significant advantage. RF struggles to capture complex spatial relationships, as reflected in its relatively high RMSE values (e.g., 99.87 in Sub-region I) and low R2 scores across sub-regions. MLP-PE, despite incorporating positional encoding, performs poorly with high errors and negative R2 in some sub-regions (e.g., Sub-region III), indicating limited generalization. In contrast, DeepTopoNet effectively leverages spatial features and multimodal data, achieving an RMSE as low as 19.38 and an SSIM of 0.94 in Sub-region I. VAE+XGBoost performs better than MLP-PE and RF but still falls short of DeepTopoNet in all sub-regions. While VAE+XGBoost benefits from variational autoencoder-based feature extraction and gradient-boosted decision trees for regression, its RMSE remains significantly higher than DeepTopoNet across all sub-regions (e.g., 108.54 in Sub-region I vs. 19.38 for DeepTopoNet).
When compared to patch-based deep learning models, including U-Net (Ronneberger et al., 2015), U-Net++ (Zhou et al., 2018), U-Net3+ (Huang et al., 2020), and Attention U-Net (Oktay et al., 2018), DeepTopoNet shows a clear improvement. Among these baselines, Attention U-Net performs the best, achieving competitive results with an RMSE of 33.33 and SSIM of 0.90 in Sub-region I. However, DeepTopoNet consistently surpasses Attention U-Net, achieving a much lower RMSE of 19.38 and a higher SSIM of 0.94 in the same sub-region. This indicates that DeepTopoNet’s combination of enriched input features, dynamic loss balancing, and robust architecture allows it to better capture spatial relationships and generalize effectively across diverse sub-regions. In summary, DeepTopoNet’s consistent outperformance across all categories of baselines demonstrates its superior capability for accurately reconstructing subglacial topography, providing a robust solution for regions with complex and sparse observational data.
@@ 5.4. Ablation Study
w/o stride, gradient (), and trend surface (t) features. Removing these components leads to a substantial decline in performance. For instance, in Sub-region IV, RMSE increases drastically to 457.06, and R2 becomes negative (e.g., -57.74), indicating a failure to capture meaningful spatial patterns. In other sub-regions, metrics such as SSIM and PSNR are lower compared to DeepTopoNet, emphasizing the importance of these features for accurate predictions.
w/o Stride, w/ gradient (), and trend surface (t) features. Retaining gradient and trend surface features while removing stride significantly improves performance compared to the first scenario. For example, in Sub-region I, RMSE reduces to 36.26, and R2 increases to 0.96, demonstrating that gradient and trend surface features are vital for capturing spatial variability. However, the performance still lags behind DeepTopoNet in certain metrics, such as RMSE in Sub-region III (27.55 vs. 32.69 for DeepTopoNet).
w/o BedMachine Loss (). Removing the BedMachine Loss () significantly degrades the model’s performance, highlighting its critical role in training DeepTopoNet. Across all sub-regions, the absence of leads to substantial increases in RMSE, particularly in Sub-region IV, where RMSE jumps to 106.00 and R2 drops to -2.16, indicating a severe loss of predictive accuracy. Additionally, SSIM and PSNR values decline across all sub-regions, reflecting a reduction in structural similarity and topographic detail retention. The poor performance suggests that plays a crucial role in constraining the predictions, ensuring that the model maintains coherence with the global topographic trends derived from BedMachine data.
DeepTopoNet. The full DeepTopoNet framework achieves the best results across most metrics and sub-regions. For example, in Sub-region I, RMSE is the lowest at 19.38, and SSIM reaches 0.94, indicating excellent structural similarity between predictions and ground truth. This highlights the effectiveness of combining stride, , t, and for robust and accurate bed topography predictions. The ablation study shows that each component contributes significantly to the performance of DeepTopoNet. While and t play a key role in improving accuracy, the inclusion of stride further enhances the model’s ability to generalize and capture spatial context, leading to superior performance across all sub-regions. Additionally, the proves to be a crucial component of the model. This confirms that is essential for balancing high-resolution radar-derived observations with global-scale topographic trends, ensuring stability and accuracy in bed topography predictions.
| Prediction/BedMachine (Morlighem et al., 2017, 2020)/Difference |
@@ 5.6. Spatial Generalization Analysis
We evaluate DeepTopoNet and representative deep learning baselines (e.g., U-Net, U-Net++, U-Net3+, and Attention U-Net) under this spatial holdout scheme. As shown in Table 6, DeepTopoNet achieves the best performance across all major evaluation metrics in Sub-region I. It produces the lowest MAE (15.90) and RMSE (25.39), while maintaining the highest (0.98), SSIM (0.91), and PSNR (33.90). Notably, even though U-Net and U-Net3+ perform moderately well, they fall short in both structural similarity and magnitude-based errors. While Attention U-Net achieves competitive performance, DeepTopoNet still outperforms it in all categories except TRI, where U-Net slightly edges out others in preserving localized ruggedness.

### v2

@@ Abstract.
Mapping Greenland’s subglacial topography is critical for projecting the future mass loss of the ice sheet and its contribution to global sea-level rise. However, the complex and sparse nature of observational data, particularly information about the bed topography under the ice sheet, significantly increases the uncertainty in model projections. Bed topography is traditionally measured by airborne ice-penetrating radars that measure the ice thickness directly underneath the aircraft, leaving data gaps of tens of kilometers in between flight lines. This study introduces a deep learning framework, DeepTopoNet, that integrates radar-derived ice thickness observations and BedMachine Greenland data through a novel dynamic loss-balancing mechanism. Among all efforts to reconstruct bed topography, BedMachine has emerged as one of the most widely used datasets, combining mass conservation principles and ice thickness measurements to generate high-resolution bed elevation estimates. The proposed loss function adaptively adjusts the weighting between radar and BedMachine’s bed, ensuring robustness in areas with limited radar coverage while leveraging the high spatial resolution of BedMachine’s bed estimates. Our approach incorporates gradient-based and trend surface features to enhance model performance and utilizes a convolutional neural network (CNN) architecture (i.e., BedTopoCNN) designed for subgrid-scale predictions. By systematically testing on the Upernavik Isstrøm) region in West Greenland, the model achieves high accuracy (MAE: 12.49 m, RMSE: 19.38 m, and R2: 0.99), outperforming baseline methods in reconstructing subglacial terrain. This work demonstrates the potential of deep learning in bridging observational gaps, providing a scalable and efficient solution to inferring subglacial topography. This framework paves the way for improved predictions of ice sheet flow and sea level rise.
@@ 1. Introduction
The Greenland Ice Sheet is a major contributor to global sea-level rise, driven by warmer atmospheric temperatures and the intrusion of warm ocean currents in fjords. Rising sea level poses a critical threat to coastal communities worldwide, increasing the risk of flooding, shoreline erosion, and habitat loss (Team et al., 2023). Estimating its contribution to sea level requires a detailed understanding of subglacial topography, which influences ice flow and basal friction (Barnes et al., 2021; Åkesson et al., 2021). However, accurately inferring the bed topography remains a significant challenge due to limited observational coverage, especially in the ice sheet interior. Traditional mapping methods, such as mass conservation models and geostatistic techniques (Morlighem et al., 2011; Bamber et al., 2001, 2013), have been instrumental but often struggle in regions of slower flow.
@@
To summarize, the contributions of this paper are threefold: (i) DeepTopoNet Framework: A novel deep learning framework integrating radar observations and BedMachine data, featuring a CNN architecture enhanced with gradient-based covariates and trend-surface features; (ii) Dynamic Loss-Balancing Mechanism: An adaptive mechanism ensuring robust performance in regions with sparse radar data by balancing contributions from multimodal data sources; and (iii) Comprehensive Evaluation: Extensive benchmarking across Greenland’s Upernavik Isstrøm sector and unseen regions, demonstrating superior performance over traditional interpolation and machine learning methods.
@@ 2. Related Work
Subglacial Bed Topography Prediction. Inferring the subglacial bed topography has been a critical focus in glaciology due to its implications for ice sheet mass balance and its contribution to global sea-level rise. Recent advancements have integrated machine learning, physics-based models, and hybrid approaches to tackle this complex problem. Cheng et al. (2024) (Cheng et al., 2024) applied Physics-Informed Neural Networks (PINNs) to model ice flow dynamics for Helheim Glacier in Greenland. The study demonstrated PINNs’ capability to infer basal friction coefficients and interpolate sparse ice thickness observations with high accuracy. The framework uniquely incorporates the conservation of momentum (i.e., the shallow-shelf/shelfy-stream approximation) (MacAyeal, 1989; Morland, 1987) as soft constraints in the loss function, improving predictions where data are sparse and enhancing the versatility of the model. Yi et al. (2023) (Yi et al., 2023) evaluated nine machine learning and statistical models, including XGBoost, Gaussian Process Regression, and hybrid models, to predict Greenland’s subglacial bed topography. The study highlighted the strong performance of XGBoost when paired with kriging-based interpolation, achieving the best RMSE and R2. Similarly, Leong and Horgan (2020) (Leong and Horgan, 2020) introduced DeepBedMap, a deep learning framework based on generative adversarial networks, to produce high-resolution bed elevation maps of Antarctica. Their approach combined multi-resolution spatial data, such as surface elevation and ice velocity, to improve the representation of subglacial terrain roughness. These studies underscore the importance of combining domain knowledge with advanced computational tools to address challenges in regions with sparse observations. This paper builds on these foundations, proposing a robust framework that integrates multimodal data, deep learning, and loss balancing to further enhance predictive performance.
@@ 4.1. Prior Physical Model (BedMachine)
BedMachine (Morlighem et al., 2017, 2020) is the most widely used subglacial topography map of Greenland. It employs a mass conservation approach in fast flowing regions, combining radar-derived ice thickness measurements, satellite velocity data, and surface mass balance reconstructions to produce a high-resolution bed topography map. The idea of mass conservation is to infer the bed elevation by ensuring conservation of mass: , where is the apparent mass balance that combines observed thickness change and surface mass balance from a regional climate model. BedMachine ensures physically realistic reconstructions in regions of fast-flowing ice, but remains highly uncertain in slower moving regions where it is based on a simple interpolation scheme that connects existing radar data. The output is a seamless topographic map with horizontal resolution up to 150 m. In this study, BedMachine data serves as a reference for the target variable, subglacial bed topography (). The proposed DeepTopoNet framework goes beyond the interpolation based approach and the reliance on mass conservation.
@@ 4.4. Loss Function
The magnitudes of and are inherently influenced by the proportion of radar-covered () and non-radar () pixels in each mini-batch. In radar-dense batches, is computed over many radar points and tends to dominate, resulting in a smaller and a larger . In radar-sparse batches, dominates, increasing and reducing . This coupling between loss magnitude and data coverage means that the weighting terms implicitly adapt to the local radar density of each batch without requiring explicit density computation or manual tuning.
@@ 5.1.1. Dataset
The dataset used in this study is obtained from the Upernavik Isstrøm region in Central West Greenland, encompassing four distinct sub-regions. Each sub-region covers a spatial extent of grid cells, representing an area with diverse ice dynamics and varying data density. The number of radar data points differs across the sub-regions, reflecting the uneven distribution of observations. The dataset includes several observed variables critical for subglacial topography inference:
@@ 5.3. Quantitative Comparison
Compared to traditional interpolation methods, such as IDW (Johnston et al., 2001) and RBF (Heryudono and Driscoll, 2010), DeepTopoNet delivers significantly improved performance. For instance, while IDW achieves moderate results in some sub-regions (e.g., SSIM of 0.87 in Sub-region II), its RMSE values remain substantially higher (e.g., 101.05 m in Sub-region I). Similarly, RBF fails to capture finer spatial details, resulting in higher errors across all metrics. In contrast, DeepTopoNet achieves considerably lower RMSE and higher SSIM, demonstrating its ability to model spatial variability more effectively than interpolation methods.
Against pixel-based machine learning methods, including RF (Sekulić et al., 2020), MLP-PE (Li et al., 2021) and VAE+XGBoost (Yi et al., 2023), DeepTopoNet exhibits a significant advantage. RF struggles to capture complex spatial relationships, as reflected in its relatively high RMSE values (e.g., 99.87 m in Sub-region I) and low R2 scores across sub-regions. MLP-PE, despite incorporating positional encoding, performs poorly with high errors and negative R2 in some sub-regions (e.g., Sub-region III), indicating limited generalization. In contrast, DeepTopoNet effectively leverages spatial features and multimodal data, achieving an RMSE as low as 19.38 m and an SSIM of 0.94 in Sub-region I. VAE+XGBoost performs better than MLP-PE and RF but still falls short of DeepTopoNet in all sub-regions. While VAE+XGBoost benefits from variational autoencoder-based feature extraction and gradient-boosted decision trees for regression, its RMSE remains significantly higher than DeepTopoNet across all sub-regions (e.g., 108.54 m in Sub-region I vs. 19.38 m for DeepTopoNet).
When compared to patch-based deep learning models, including U-Net (Ronneberger et al., 2015), U-Net++ (Zhou et al., 2018), U-Net3+ (Huang et al., 2020), and Attention U-Net (Oktay et al., 2018), DeepTopoNet shows a clear improvement. Among these baselines, Attention U-Net performs the best, achieving competitive results with an RMSE of 33.33 m and SSIM of 0.90 in Sub-region I. However, DeepTopoNet consistently surpasses Attention U-Net, achieving a much lower RMSE of 19.38 m and a higher SSIM of 0.94 in the same sub-region. This indicates that DeepTopoNet’s combination of enriched input features, dynamic loss balancing, and robust architecture allows it to better capture spatial relationships and generalize effectively across diverse sub-regions. In summary, DeepTopoNet’s consistent outperformance across all categories of baselines demonstrates its superior capability for accurately reconstructing subglacial topography, providing a robust solution for regions with complex and sparse observational data.
@@ 5.4. Ablation Study
w/o stride, gradient (), and trend surface (t) features. Removing these components leads to a substantial decline in performance. For instance, in Sub-region IV, RMSE increases drastically to 457.06 m, and R2 becomes negative (e.g., -57.74), indicating a failure to capture meaningful spatial patterns. In other sub-regions, metrics such as SSIM and PSNR are lower compared to DeepTopoNet, emphasizing the importance of these features for accurate predictions.
w/o Stride, w/ gradient (), and trend surface (t) features. Retaining gradient and trend surface features while removing stride significantly improves performance compared to the first scenario. For example, in Sub-region I, RMSE reduces to 36.26 m, and R2 increases to 0.96, demonstrating that gradient and trend surface features are vital for capturing spatial variability. However, the performance still lags behind DeepTopoNet in certain metrics, such as RMSE in Sub-region III (27.55 m vs. 32.69 m for DeepTopoNet).
w/o BedMachine Loss (). Removing the BedMachine Loss () significantly degrades the model’s performance, highlighting its critical role in training DeepTopoNet. Across all sub-regions, the absence of leads to substantial increases in RMSE, particularly in Sub-region IV, where RMSE jumps to 106.00 m and R2 drops to -2.16, indicating a severe loss of predictive accuracy. Additionally, SSIM and PSNR values decline across all sub-regions, reflecting a reduction in structural similarity and topographic detail retention. The poor performance suggests that plays a crucial role in constraining the predictions, ensuring that the model maintains coherence with the global topographic trends derived from BedMachine data.
DeepTopoNet. The full DeepTopoNet framework achieves the best results across most metrics and sub-regions. For example, in Sub-region I, RMSE is the lowest at 19.38 m, and SSIM reaches 0.94, indicating excellent structural similarity between predictions and ground truth. This highlights the effectiveness of combining stride, , t, and for robust and accurate bed topography predictions. The ablation study shows that each component contributes significantly to the performance of DeepTopoNet. While and t play a key role in improving accuracy, the inclusion of stride further enhances the model’s ability to generalize and capture spatial context, leading to superior performance across all sub-regions. Additionally, the proves to be a crucial component of the model. This confirms that is essential for balancing high-resolution radar-derived observations with global-scale topographic trends, ensuring stability and accuracy in bed topography predictions.
| Prediction — BedMachine (Morlighem et al., 2017, 2020) — Difference |
@@ 5.6. Spatial Generalization Analysis
We evaluate DeepTopoNet and representative deep learning baselines (e.g., U-Net, U-Net++, U-Net3+, and Attention U-Net) under this spatial holdout scheme. As shown in Table 6, DeepTopoNet achieves the best performance across all major evaluation metrics in Sub-region I. It produces the lowest MAE (15.90 m) and RMSE (25.39 m), while maintaining the highest R2 (0.98), SSIM (0.91), and PSNR (33.90). Notably, even though U-Net and U-Net3+ perform moderately well, they fall short in both structural similarity and magnitude-based errors. While Attention U-Net achieves competitive performance, DeepTopoNet still outperforms it in all categories except TRI, where U-Net slightly edges out others in preserving localized ruggedness.
@@ Appendix A Code Availability
Appendix B Additional Results for Sub-region II, III, and IV
To further evaluate the performance of the proposed framework and baseline methods, we present additional results for Sub-regions II, III, and IV (see Table 7). These regions provide diverse challenges due to variations in radar data density, terrain complexity, and topographic features.
| Sub-region II — Sub-region III — Sub-region IV |
| IDW (Johnston et al., 2001) |
| RBF (Heryudono and Driscoll, 2010) |
| RF (Sekulić et al., 2020) |
| MLP-PE (Li et al., 2021) |
| VAE+XGBoost (Yi et al., 2023) |
| U-Net (Ronneberger et al., 2015) |